From: STOSC::SMM "On-Duty SMM x5074" 11-JAN-1993 21:11:04.23 To: @[SMM.DIS]HST_STAT CC: Subj: HUBBLE SPACE TELESCOPE DAILY REPORT #800 PERIOD COVERED: 0800 EST 01/08/93 - 0800 EST 01/11/93 Daily Status Report as of 011/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed GHRS/2 4217 (Absorption Cross sections of Nearby Galaxies) The High Resolution Spectrograph was used to make a single spectral observation of quasar 1341+258. The metal absorption lines in the quasar are thought to arise from matter associated with intervening galaxies. The observations was received and displayed. The spectrum was weak, which was expected because the target was very faint. 1.2 Completed PC 4084 (Globular Cluster Cores, GTO Augmentation) The Planetary Camera was used to take eight U-band images of the core of globular cluster NGC-5272 (aka, M-3 located in Canus Venatici). All of the images were received and displayed. They all were reported to look okay. 1.3 Completed FOC/48 4740 (Intensifier HV Test) The high voltage was brought up on the Faint Object Camera (f/48) for the first time since the intensifier high voltage anomaly of Sept. 17,1992. The high voltage was turned on just before LOS. At the AOS the high voltage was off. It was not clear what the problem was. An HSTAR was written. Seven observations were received and displayed. They all were blank because the high voltage went off shortly after it came on. (See Operations Note 2.4) 1.4 Completed PC 4140 (Rapid Internal Monitor) The Planetary Camera was used to take two internal flat images to be used to check for new forms of contamination. Both images were received and displayed in the OSS. They both were reported to look okay. 1.5 Completed Two Sets of HSP/UV2/POL 3234 (X-Ray Binaries) The High Speed Photometer was used to make make both photometric and polarimetric observations of the x-ray binary star HD77581 (aka, Vela XR-1). 18 observations were taken on the first set. They ware all received and displayed. Most of them looked okay, four of them had short term fluctuations significantly higher than the background noise level. This was apparently due to spacecraft jitter associated with a terminator crossing. The second set of 18 observations were also received and displayed. The first two images were target acquisition images, and showed the target well centered. The following observations looked good, but two images were filed as bad because they had some short term fluctuations similar to those in the first set. 1.6 Completed HSP/UV2 3255 (Optical Variability Associated with Black Holes) The High Speed Photometer was used to try and obtain high rate ultraviolet photometry observations of x-ray binary star A0620-00 whose companion is a suspected black hole. The target acquisition images were blank. The time series data had low counts. 1.7 Completed WFC 1306 (GTO Low Latitude Parallel Program) The Wide Field Camera was used to make some low latitude observations in parallel with the prime science observations. All observations were received and displayed in the OSS. They were all reported to be nominal. 1.8 Completed FOS/RD 4212 (Internal Wave Length Calibration) The Faint Object Spectrograph was used to obtain wave length calibration images using every dispenser-detector combination. All observations used internal sources. The observations were received and displayed, and all were reported to be okay. 1.9 Completed Two Sets of FOC/96 4107 (Search for QSOs Suitable for Subsequent Observations) The Faint Object Camera was used to try and obtain images of three faint quasars (i.e., 1013+00, 0307+0222, and 0058-276B) to see if they had sufficient ultraviolet flux for future observations. The observations were received and displayed, however, they were all reported to be blank. 1.10 Completed FOC/96 3801 (Search to QSOs Suitable for Observation) The Faint Object Camera Was used to make one ultraviolet observation of faint high red shift quasar 0938+199. The observations was received and displayed. The image was blank. 1.11 Completed FOC/96 4075 (Magellanic Cloud Planetary Nebulae) The Faint Object Camera was used to make a single image of the planetary nebula SMC-N67 located in the Small Magellanic Cloud. The image will be used to measure its diameter. The image was reported to look good. 1.12 Completed FOS/BL 4110 (The Physics of Massive O Stars) The Faint Object Spectrograph was used to make six observations of massive OB class star SK67D211 which is located in the Large Magellanic Cloud. Ultraviolet spectral observations were taken from 1150 to 2300 A0. All six observations were reported to look okay. 1.13 Completed WFC 1297 (GTO High Latitude Parallel Program) The Wide Field Camera was used to take a series of four high latitude images in parallel with the prime science observations. All four of the images were reported to look okay. 1.14 Completed WFC 4105 (HST Medium Deep Survey: High Latitude) The Wide Field Camera was used to take a single image as part of a basic medium-deep sky survey taken in parallel with the primary science observations. The observations was reported to look okay. 1.15 Uplinked Updated Ephemeris Tables, Per MOSES 2871 1.16 Completed WFC-PC 4109 (Rapid Internal Monitor) Both the Wide Field Camera and the Planetary Camera were used to make internal flat images looking for new forms of contamination (i.e., daisies and measles). Both images were inspected in the OSS, and reported to be okay. 1.17 Completed FGS-PC 4765 (Overlapping Plate Field Distortion) Both the Planetary Camera and Fine Guidance Sensor #3 were used to measure the optical field angle distortion (OFAD). The Planetary Camera observations were displayed, and there was at least one star image in each observation. The astrometry observations were monitored, and no problems were reported in the telemetry. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 29 Successful : 29 Scheduled Reacquisitions : 13 Successful : 13 Losses of Lock : 1 (LOS at EON 011/0826Z) 2.2 FHST Updates: Scheduled : 27 Successful : 27 2.3 Significant HSTARs: None 2.4 Operations Note: On Day 008/1926Z The Faint Object Camera (f/48) was powered up to fulfill a CARD restriction that requires that the FOC Camera P/S to be cycled every 30 days. This restriction has been waived once due to the IEHT anomaly that occurred on Sept. of 1992, so it had been a little over two months since the last cycling event. The IEHT has been powered up successfully two times since the original failure. The turn on procedure that was attempted on Day 008 employed the modified ramp up procedure that was used in previous engineering tests and appeared to be proceeding nominally as the spacecraft entered LOS. But at AOS the IEHT telemetry indicated it was powered off. The implication is that the PDA H/W current limit was tripped, powering off the IEHT. ESA and STScI engineers are investigating the anomaly. (See Note 1.3) 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete FGS-PC 4765 (Overlapping Plate Field Distortion) 3.2 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.3 Complete HSP/POL/UV2 3234 (X-Ray Binaries) 3.4 Complete Four Sets of WFC 4713 (Internal Delta Flats) 3.5 Complete FOS/BL 3573 (The reflected Light Model for Seyfert) 3.6 Complete FOS/BL 4110 (The Physics of Massive O-Stars) 3.7 Complete WFC 4105 (HST Medium Deep Survey: High Latitude) 3.8 Complete FGS 4031 (Search for Planetary Companions to Low Mass) 3.9 Complete PC 3848 (Separate Light curves of Pluto and Charon) 3.10 Complete PC 4714 (Internal Delta Flats) 3.11 Complete FOS/RD 3840 (Abundance and Time Evolution of Carbon) 3.12 Initiate WFC 3917 (HST Medium Deep Survey, Part 1) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time -END-